OpenStudy (anonymous):

can someone please answer this to me?? can you tell me in an equation how to find 20% of 1 liter?????

OpenStudy (anonymous):

20 percent means = 20/100 so 1 times (20/100) =

OpenStudy (anonymous):

never mind not of 1 liter, its more like 1 liter plus 20%. but what is the total? can u help me set up an equation for tht??

OpenStudy (anonymous):

with only one variable and that one variable would be the total

OpenStudy (anonymous):

well when you find what 20 percent of 1 liter is you can add that to 1 liter to find the total. Equation would be 1 liter + 20% of 1 liter = total 1 + (20/100) = total total= ?

OpenStudy (anonymous):

let x= total then x= 1 + (20/100) same as x= 1+ 0.2

OpenStudy (anonymous):

1.2+x then??

OpenStudy (anonymous):

=x

OpenStudy (anonymous):

not add

OpenStudy (anonymous):

yes, x= 1.2 :)

OpenStudy (anonymous):

thank you so much. it all makes a lot of sense to me now!!!

OpenStudy (anonymous):

You're welcomed! :)